solve the equation:
2/t = t/3t - 4

Answer by TimothyLamb(4379)

2/t = t/3t - 4
2/t = (t/t)(1/3) - 4
2/t = (1/3) - 4
2 = ( (1/3) - 4 )t
t = 2/( (1/3) - 4 )
t = -0.545454545455
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php